# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 9.4 Line 11: A 5-vertex with fewer than three 4-neighbors
We suppose we are in Branch5-2 here. Let $U$ be the non-empty set of 5-vertices which have fewer than three 4-neighbors. For any vertex $u\in U$, define the induced neighborhood $H_{u}=G[N(u)]$. Note that any 4-vertex is isolated in $H_{u}$, as otherwise we would apply Line 10 to it with its 5-neighbor $u$. The following theorem characterizes the possible neighborhood structure (see Figure 7).
@@
There is some vertex $u\in U$ and an enumeration of its neighbors $v_{1},\dots,v_{5}$ to satisfy one of the following three conditions:
@@
    $H_{u}$ is a matching and $v_{1}\sim v_{2},\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=5,\deg(v_{5})=4,\deg_{G-v_{3}}(v_{4})=4$ and $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$.
@@
    $v_{1}\sim v_{2},v_{3}\sim v_{4},\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=\deg(v_{4})=5$. Furthermore, there are 5-vertices $x_{13},x_{24},y_{2},y_{4}$ and 4-vertices $z_{2},z_{4}$, all distinct from each other and from $N[u]$, such that $v_{2},x_{24},y_{2}$ and $v_{4},x_{24},y_{4}$ are both triangles in $G$ and $x_{2}\sim x_{24}\sim x_{4}$ and $x_{13}\not\sim\{x_{24},v_{2},v_{4}\}$ and $v_{2}\sim z_{2},v_{4}\sim z_{4}$.
@@
If $H_{u}$ has no edges, it falls into the first category (choose $v_{1},v_{2},v_{3}$ to be 5-neighbors of $u$).
Suppose $H_{u}$ is not a matching, i.e. there is a neighbor $v_{1}$ with $\operatorname{codeg}(v_{1},u)\geq 2$. Let $v_{2},v_{3}$ be two shared neighbors of $v_{1},u$ and let $v_{4},v_{5}$ be any other neighbors of $u$. Note that $G-\{v_{4},v_{5}\}$ has a kite. Since 4-vertices are isolated in $H_{u}$, the only possible structure where we would not use the branching rule at Line 3 must be that $\deg(v_{1})=5,\deg(v_{4})=\deg(v_{5})=4$; since 4-vertices are isolated in $H_{u}$ necessarily $\{v_{1},v_{2},v_{3}\}\not\sim\{v_{4},v_{5}\}$. This gives Case III.
Suppose $H_{u}$ is a matching with exactly one edge $(v_{1},v_{2})$, and an additional 5-vertex $v_{3}$. If it has any other 5-neighbor $t$, then $R(v_{3},t)\geq 1$ (due to 3-triangle $u,v_{1},v_{2}$) and we would branch at Line 9. So $u$ must have two 4-neighbors $v_{4},v_{5}$. Then Observation 54 implies $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$. This gives Case II.
@@
Consider some $u\in U$, where $H_{u}$ has edges $(v_{1},v_{2}),(v_{3},v_{4})$ and has a 4-vertex $v_{5}$. By Observation 52 we automatically have $\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=1$. We claim that $\operatorname{codeg}(v_{3},v_{4})=1$. For, suppose $v_{3},v_{4}$ share neighbor $t\neq u$; since $t$ has two 5-neighbors necessarily $\deg(t)=5$. Then $v_{3}\in U$ since it has three 5-neighbors $u,v_{4},t$, but $\deg_{H_{v_{3}}}(v_{4})\geq 2$, which contradicts our assumption on structure of $U$. By completely symmetric arguments also $\operatorname{codeg}(v_{1},v_{2})=1$.
Now consider forming a bipartite graph $J$ with left-nodes $v_{1},v_{2}$ and right-nodes $v_{3},v_{4}$, with an edge $(v_{i},v_{j})$ if $\operatorname{codeg}_{G}(v_{i},v_{j})=2$. If any node has degree zero in $J$, we may suppose without loss of generality that $\deg_{J}(v_{4})=0$ in which case $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=1$ giving us Case II.
Otherwise, if every node in $J$ has at least one neighbor, then by Hall’s theorem $J$ has a perfect matching; by swapping vertices as needed we assume this matching is $(v_{1},v_{3}),(v_{2},v_{4})$, that is, $v_{1},v_{3}$ share a neighbor $x_{13}\neq u$ and $v_{2},v_{4}$ share a neighbor $x_{24}\neq u$. Here, $\deg(x_{13})=\deg(x_{24})=5$ since they each have two 5-neighbors. Clearly $u\not\sim\{x_{13},x_{24}\}$. By Observation 52, we have $\operatorname{codeg}(x_{13},u)\leq 2$ and $\operatorname{codeg}(x_{24},u)\leq 2$, so $x_{13}\not\sim\{v_{2},v_{4}\}$ and $x_{24}\not\sim\{v_{1},v_{3}\}$. In particular, $x_{13}\neq x_{24}$. Also, $x_{13}\not\sim x_{24}$, as otherwise $x_{13}$ would have three independent 5-neighbors $x_{24},v_{1},v_{3}$, contradicting our assumption of structure of $U$.
Now $v_{2}\in U$ as it has three 5-neighbors $v_{1},u,x_{24}$ and similarly $v_{4}\in U$ as it has three 5-neighbors $v_{3},u,x_{24}$. By our hypothesis on the structure of $U$, it must be that $v_{2}$ has an edge $(x_{24},y_{2})$ in its neighborhood and has a 4-neighbor $z_{2}$, and likewise $v_{4}$ has an edge $(x_{24},y_{4})$ in its neighborhood and has a 4-neighbor $z_{4}$. By Observation 52, we have $\operatorname{codeg}(x_{24},u)\leq 2$, and so $y_{2},y_{4}\notin N[u]$. Also $y_{2}\neq y_{4}$, as otherwise the vertex $y=y_{2}=y_{4}$ would have $\operatorname{codeg}(v_{2},v_{4})\geq 3$, sharing neighbors $u,x_{24},y$. Since $x_{13}\not\sim x_{24}$, it must be that $y_{2},y_{4}$ are distinct from $x_{13}$. Likewise, we have $z_{2}\neq z_{4}$ since otherwise the vertex $z=z_{2}=z_{4}$ would have two 5-neighbors $v_{2},v_{4}$. This gives Case IV. ∎
@@
Case I. Split on $v_{1},v_{2},v_{3}$, getting subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}$. By Observation 52, these have drops $(2,5),(2.5,6),(3,7),(1,3)$ directly. In $G_{4}$, we apply Proposition 30 to 2-vertex $u$, where by Observation 52 and Observation 54 we have $r=\deg_{G_{4}}(v_{4})+\deg_{G_{4}}(v_{5})-\operatorname{codeg}_{G_{4}}(v_{4},v_{5})-1\geq 6$ irrespective of the degrees of $v_{4},v_{5}$. We need:
@@
Case IIa: $\boldsymbol{\operatorname{shad}(v_{3},N[v_{1},v_{4}])\geq 1}$. Split on $v_{3},v_{5}$ and apply (A2) to vertex $v_{1}$ in $G-\{v_{3},v_{5}\}$. This generates subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{5}],G_{3}=G-\{v_{3},v_{5}\}-N[v_{1},v_{4}],G_{4}=G-\{v_{1},v_{3},v_{5}\}$. Since $G$ is in 5BP, and $\deg_{G}(v_{3})=5,\deg_{G}(v_{5})=4,v_{3}\not\sim v_{5}$, subproblems $G_{1},G_{2}$ both have drop $(2,5)$ directly. Case IIa gives $\operatorname{minsurp}(G_{3})\geq 0$; since $\operatorname{codeg}(v_{1},v_{4})=1$, it has drop $(4,10)$. Subproblem $G_{4}$ has drop $(1.5,3)$ directly, and we also apply Proposition 30 to 2-vertex $u$ in $G_{4}$, noting that $\deg_{G_{4}}(v_{2})+\deg_{G_{4}}(v_{4})-1-\operatorname{codeg}_{G_{4}}(v_{2},v_{4})=4+4-1-1=6$.
@@
Case IIb: $\boldsymbol{\operatorname{shad}(v_{3},N[v_{1},v_{4}])\leq 0}$. Split on $v_{3},v_{4}$, then apply (A2) to vertex $v_{1}$ in $G-\{v_{3},v_{4}\}$. We get subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{4}],G_{3}=G-\{v_{3},v_{4}\}-N[v_{1},v_{5}],G_{4}=G-\{v_{1},v_{3},v_{4}\}$. Since $G$ is in 5BP, subproblems $G_{1},G_{2}$ both have drop $(2,5)$.
In subproblem $G_{2}$, consider vertex $v_{1}$; it has degree 4 since $\operatorname{codeg}(v_{1},v_{4})=1$ and $v_{1}\not\sim v_{3}$. Case IIb gives $\operatorname{shad}_{G_{2}}(v_{1})=\operatorname{shad}(v_{3},N[v_{1},v_{4}])\leq 0$, that is, $v_{1}$ is a blocked 4-vertex in $G_{2}$. Also, $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[v_{5}])-1\geq 0$. So, either $S(G_{2})\geq 1$ or we can apply Proposition 25 to $G_{2}$.
@@
Case IV: Split on the independent 5-vertices $u,x_{13},x_{24}$, generating subproblems $G_{1}=G-N[u],G_{2}=G-u-N[x_{13}],G_{3}=G-\{u,x_{13}\}-N[x_{24}],G_{4}=G-\{x_{14},x_{23},u\}$. Note that $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{24}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{14})-2\geq 0$. So these have drops $(2,5),(2.5,6),(3,7),(1.5,3)$ directly.
Consider graph $G_{2}$. By Observation 52, we have $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[x_{13}])-1\geq 1$, in particular $\operatorname{mindeg}(G_{2})\geq 2$. Also note that $v_{4}\not\sim v_{4}$ and both $z_{2},z_{4}$ remain in the graph $G_{2}$, since $z_{2},z_{4}$ are 4-vertices which already have a 5-neighbors $v_{2}$ or $v_{4}$ respectively.
We claim that $S(G_{2})\geq 2$; there are a few different cases depending on $|N(x_{13})\cap\{y_{2},y_{4}\}|$. First, suppose $x_{13}\sim y_{2}$ and $x_{13}\sim y_{4}$; then, in the graph $G_{2}$, vertices $v_{2}$ and $v_{4}$ have degree two (losing neighbors $v_{1},u,y_{2}$ and $v_{3},u,y_{4}$ respectively). So by Observation 11 we immediately have $S(G_{2})\geq 2$. Next, suppose $|N(x_{13})\cap\{y_{2},y_{4}\}|=1$, say without loss of generality $x_{13}\sim y_{2},x_{13}\not\sim y_{4}$. In this case, we apply (P2) to the 2-vertex $v_{2}$ in $G_{2}$; in the resulting graph $G_{2}^{\prime}$, the vertices $z_{2},x_{24}$ get merged into a new vertex $z^{\prime}$. In $G_{2}^{\prime}$, the vertex $v_{4}$ has neighbors $z_{4},z^{\prime},y_{4}$ where $z^{\prime}\sim y_{4}$ in $G_{2}^{\prime}$ since $x_{24}\sim y_{4}$ in $G_{2}$. So we can apply (P3) to 3-triangle $v_{2}^{\prime}$, yielding $S(G_{2}^{\prime})\geq 1$ and hence $S(G_{2})\geq 2$. Finally suppose $x_{13}\not\sim\{y_{2},y_{4}\}$; then in $G_{2}$, the vertex $v_{2}$ has degree 3 with a triangle $v_{1},y_{2}$. We apply (P3) to $v_{2}$; let $G_{2}^{\prime}$ denote the resulting graph, where we remove $z_{2}$ and add edges from $v_{1},y_{2}$ to the neighbors of $z_{2}$. Since $v_{4}\not\sim v_{2}$, the degree of $v_{4}$ does not change in $G_{2}^{\prime}$, and in particular $v_{4}$ still has degree 3 with a triangle $v_{3},y_{4}$. So $S(G_{2}^{\prime})\geq 1$ and $S(G_{2})\geq 2$.
Consider graph $G_{3}$; here, vertex $v_{1}$ loses neighbors $u,x_{13},v_{2}$ and likewise $v_{3}$ loses neighbors $u,x_{13},v_{4}$). Moreover $\operatorname{minsurp}(G_{3})\geq\operatorname{minsurp}(G_{24})-2\geq 0$. Thus, by Observation 11 we have $S(G_{3})\geq 2$.

### v7

@@ 7.4 Line 11: A 5-vertex with fewer than three 4-neighbors
We suppose we are in Branch5-2. Let $U$ be the non-empty set of 5-vertices with fewer than three 4-neighbors. For any vertex $u\in U$, define the induced neighborhood $H_{u}=G[N(u)]$. Note that any 4-vertex is isolated in $H_{u}$, as otherwise we would apply Line 10 to it with its 5-neighbor $u$. The following theorem characterizes the possible neighborhood structure (see Figure 10).
@@
There is a vertex $u\in U$ and an enumeration of its neighbors $v_{1},\dots,v_{5}$ to satisfy one of the following three conditions:
@@
    $H_{u}$ is a matching with an edge $(v_{1},v_{2})$ and possibly an edge $(v_{3},v_{4})$, where $\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=5,\deg(v_{5})=4,\deg_{G-v_{3}}(v_{4})=4$ and $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$.
@@
    $H_{u}$ is a matching with edges $(v_{1},v_{2})$ and $(v_{3},v_{4})$, where $\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=\deg(v_{4})=5$. Furthermore, there are 5-vertices $x_{13},x_{24},y_{2},y_{4}$ and 4-vertices $z_{2},z_{4}$, all distinct from each other and from $N[u]$, such that $v_{2},x_{24},y_{2}$ and $v_{4},x_{24},y_{4}$ are both triangles in $G$ and $x_{2}\sim x_{24}\sim x_{4}$ and $x_{13}\not\sim\{x_{24},v_{2},v_{4}\}$ and $v_{2}\sim z_{2},v_{4}\sim z_{4}$.
@@
If $H_{u}$ has no edges, it falls into the Case I (choose $v_{1},v_{2},v_{3}$ to be 5-neighbors of $u$).
Suppose $H_{u}$ is not a matching. Let $v_{1}$ be a neighbor which shares two other neighbors $v_{2},v_{3}$ with $u$. Let $v_{4},v_{5}$ be the other neighbors of $u$. Note that $G-\{v_{4},v_{5}\}$ has a kite. Since 4-vertices are isolated in $H_{u}$, the only possible structure where we would not use the branching rule at Line 3 is $\deg(v_{1})=5,\deg(v_{4})=\deg(v_{5})=4$; since 4-vertices are isolated in $H_{u}$ necessarily $\{v_{1},v_{2},v_{3}\}\not\sim\{v_{4},v_{5}\}$. This is Case III.
Suppose $H_{u}$ is a matching with exactly one edge $(v_{1},v_{2})$, and an additional 5-vertex $v_{3}$. If $u$ has any other 5-neighbor $t$, then $S(v_{3},t)\geq 1$ (due to 3-triangle $u,v_{1},v_{2}$) and we would branch at Line 9. So $u$ must have two 4-neighbors $v_{4},v_{5}$. Then Observation 55 implies $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$. This is Case II.
@@
Consider some $u\in U$, where $H_{u}$ has edges $(v_{1},v_{2}),(v_{3},v_{4})$ and has a 4-vertex $v_{5}$. By Observation 53 we automatically have $\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=1$. We claim that $\operatorname{codeg}(v_{3},v_{4})=1$. For, suppose $v_{3},v_{4}$ share neighbor $t\neq u$; since $t$ has two 5-neighbors necessarily $\deg(t)=5$. Then $v_{3}\in U$ since it has three 5-neighbors $u,v_{4},t$, but $\deg_{H_{v_{3}}}(v_{4})\geq 2$, which contradicts our assumption on structure of $U$. By completely symmetric arguments also $\operatorname{codeg}(v_{1},v_{2})=1$.
Now consider forming a bipartite graph $J$ with left-nodes $v_{1},v_{2}$ and right-nodes $v_{3},v_{4}$, with an edge $(v_{i},v_{j})$ if $\operatorname{codeg}_{G}(v_{i},v_{j})=2$. If any node has degree zero in $J$, we may suppose without loss of generality it is $v_{4}$, in which case case $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=1$ giving us Case II.
Otherwise, if every node in $J$ has at least one neighbor, then by Hall’s theorem $J$ has a perfect matching; by swapping vertices as needed we assume this matching is $(v_{1},v_{3}),(v_{2},v_{4})$, that is, $v_{1},v_{3}$ share a neighbor $x_{13}\neq u$ and $v_{2},v_{4}$ share a neighbor $x_{24}\neq u$. Here, $\deg(x_{13})=\deg(x_{24})=5$ since they each have two 5-neighbors. Clearly $u\not\sim\{x_{13},x_{24}\}$. By Observation 53, we have $\operatorname{codeg}(x_{13},u)\leq 2$ and $\operatorname{codeg}(x_{24},u)\leq 2$, so $x_{13}\not\sim\{v_{2},v_{4}\}$ and $x_{24}\not\sim\{v_{1},v_{3}\}$. In particular, $x_{13}\neq x_{24}$. Also, $x_{13}\not\sim x_{24}$, as otherwise $x_{13}$ would have three independent 5-neighbors $x_{24},v_{1},v_{3}$, contradicting our assumption of structure of $U$.
Now $v_{2}\in U$ as it has three 5-neighbors $v_{1},u,x_{24}$ and similarly $v_{4}\in U$ as it has three 5-neighbors $v_{3},u,x_{24}$. By our hypothesis on the structure of $U$, it must be that $v_{2}$ has an edge $(x_{24},y_{2})$ in its neighborhood and has a 4-neighbor $z_{2}$, and likewise $v_{4}$ has an edge $(x_{24},y_{4})$ in its neighborhood and has a 4-neighbor $z_{4}$. By Observation 53, we have $\operatorname{codeg}(x_{24},u)\leq 2$, and so $y_{2},y_{4}\notin N[u]$. Also $y_{2}\neq y_{4}$, as otherwise the vertex $y=y_{2}=y_{4}$ would have $\operatorname{codeg}(v_{2},v_{4})\geq 3$, sharing neighbors $u,x_{24},y$. Since $x_{13}\not\sim x_{24}$, it must be that $y_{2},y_{4}$ are distinct from $x_{13}$. Likewise, we have $z_{2}\neq z_{4}$ since otherwise the vertex $z=z_{2}=z_{4}$ would have two 5-neighbors $v_{2},v_{4}$. This gives Case IV. ∎
@@
Case I. Split on $v_{1},v_{2},v_{3}$, getting subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}$. By Observation 53, these have drops $(2,5),(2.5,6),(3,7),(1,3)$ directly. In $G_{4}$, we apply Proposition 30 to 2-vertex $u$, where by Observation 53 and Observation 55 we have $r=\deg_{G_{4}}(v_{4})+\deg_{G_{4}}(v_{5})-\operatorname{codeg}_{G_{4}}(v_{4},v_{5})-1\geq 6$ irrespective of the degrees of $v_{4},v_{5}$. We need:
@@
Case IIa: $\boldsymbol{\operatorname{shad}(v_{3},N[v_{1},v_{4}])\geq 1}$. Split on $v_{3},v_{5}$ and apply (B-Fun2) to vertex $v_{1}$ in $G-\{v_{3},v_{5}\}$. This generates subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{5}],G_{3}=G-\{v_{3},v_{5}\}-N[v_{1},v_{4}],G_{4}=G-\{v_{1},v_{3},v_{5}\}$. Since we are after Line 1, and $\deg_{G}(v_{3})=5,\deg_{G}(v_{5})=4,v_{3}\not\sim v_{5}$, the subproblems $G_{1},G_{2}$ both have drop $(2,5)$ directly. Since $\operatorname{codeg}(v_{1},v_{4})=1$, and since we are in Case IIa with $\operatorname{minsurp}(G_{3})\geq 0$, the subproblem $G_{3}$ has drop $(4,10)$. Subproblem $G_{4}$ has drop $(1.5,3)$ directly, and we also apply Proposition 30 to 2-vertex $u$ in $G_{4}$, noting that $\deg_{G_{4}}(v_{2})+\deg_{G_{4}}(v_{4})-1-\operatorname{codeg}_{G_{4}}(v_{2},v_{4})=4+4-1-1=6$.
@@
Case IIb: $\boldsymbol{\operatorname{shad}(v_{3},N[v_{1},v_{4}])\leq 0}$. Split on $v_{3},v_{4}$, then apply (B-Fun2) to vertex $v_{1}$ in $G-\{v_{3},v_{4}\}$. We get subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{4}],G_{3}=G-\{v_{3},v_{4}\}-N[v_{1},v_{5}],G_{4}=G-\{v_{1},v_{3},v_{4}\}$. Subproblems $G_{1},G_{2}$ both have drop $(2,5)$.
In subproblem $G_{2}$, consider vertex $v_{1}$; it has degree 4 since $\operatorname{codeg}(v_{1},v_{4})=1$ and $v_{1}\not\sim v_{3}$. Case IIb gives $\operatorname{shad}_{G_{2}}(v_{1})=\operatorname{shad}(v_{3},N[v_{1},v_{4}])\leq 0$, that is, $v_{1}$ is a blocked 4-vertex in $G_{2}$. Also, $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[v_{4}])-1\geq 0$. So, either $S(G_{2})\geq 1$ or we can apply Proposition 25 to $G_{2}$.
@@
Case IV: Split on the independent 5-vertices $u,x_{13},x_{24}$, generating subproblems $G_{1}=G-N[u],G_{2}=G-u-N[x_{13}],G_{3}=G-\{u,x_{13}\}-N[x_{24}],G_{4}=G-\{x_{14},x_{23},u\}$. By Observation 53, we have $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[x_{13}])-1\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{24}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{14})-2\geq 0$. So these have drops $(2,5),(2.5,6),(3,7),(1.5,3)$ directly.
Note that both $z_{2},z_{4}$ remain in the graph $G_{2}$, since $z_{2},z_{4}$ are 4-vertices which already have a 5-neighbors $v_{2}$ or $v_{4}$ respectively. We claim that $S(G_{2})\geq 2$; there are a few different cases depending on $|N(x_{13})\cap\{y_{2},y_{4}\}|$. First, suppose $x_{13}\sim y_{2}$ and $x_{13}\sim y_{4}$; then, in the graph $G_{2}$, vertices $v_{2}$ and $v_{4}$ have degree two (losing neighbors $v_{1},u,y_{2}$ and $v_{3},u,y_{4}$ respectively). So by Proposition 10 we immediately have $S(G_{2})\geq 2$. Next, suppose $|N(x_{13})\cap\{y_{2},y_{4}\}|=1$, say without loss of generality $x_{13}\sim y_{2},x_{13}\not\sim y_{4}$. In this case, we apply (P2) to the 2-vertex $v_{2}$ in $G_{2}$; in the resulting graph $G_{2}^{\prime}$, the vertices $z_{2},x_{24}$ get merged into a new vertex $z^{\prime}$. We then have $N_{G_{2}^{\prime}}(v_{4})=\{z_{4},z^{\prime},y_{4}\}$ where $z^{\prime}\sim y_{4}$, since $x_{24}\sim y_{4}$ in $G_{2}$. So we can apply (P3) to the 3-triangle on $v_{4}$, yielding $S(G_{2}^{\prime})\geq 1$ and hence $S(G_{2})\geq 2$. Finally suppose $x_{13}\not\sim\{y_{2},y_{4}\}$; then in $G_{2}$, the vertex $v_{2}$ has degree 3 with a triangle $v_{1},y_{2}$. We apply (P3) to $v_{2}$; let $G_{2}^{\prime}$ denote the resulting graph, where we remove $z_{2}$ and add edges from $v_{1},y_{2}$ to the neighbors of $z_{2}$. Since $v_{4}\not\sim v_{2}$, the degree of $v_{4}$ does not change in $G_{2}^{\prime}$, and in particular $v_{4}$ still has degree 3 with a triangle $v_{3},y_{4}$. So $S(G_{2}^{\prime})\geq 1$ and $S(G_{2})\geq 2$.
Consider graph $G_{3}$; here, vertex $v_{1}$ loses neighbors $u,x_{13},v_{2}$ and likewise $v_{3}$ loses neighbors $u,x_{13},v_{4}$. Thus, by Proposition 10 we have $S(G_{3})\geq 2$.